OpenStudy (anonymous):

4a^2 -b^2 -8bc -16c^2 factor

OpenStudy (anonymous):

-8c(2c+b)(2a+b)(2a-b)

OpenStudy (anonymous):

4a^2 - (b^2 + 8bc + 16c^2) = 4a^2 - (b + 4c)^2 = (2a - (b + 4c))(2a + (b + 4c)) = (2a - b - 4c)(2a + b + 4c)